yoboijarvis:

More help plsss ._. ss below vv Math question: For doing such an excellent job, the waitress gets a 10% raise on her hourly wage. What's her new wage per hour?

yoboijarvis:

yoboijarvis:

@XioGonz

yoboijarvis:

@AZ

XioGonz:

AZ will help you (:

AZ:

First let's calculate her hourly wage She makes $180 dollars in 20 hours How much money does she make in 1 hour? Just divide 180 / 20 = ??

yoboijarvis:

9

AZ:

So her hourly rate is $9 per hour and she gets a 10% raise so that means we're INCREASING her wage by 10% what is 10% of 9? basically 10% * 9 = ??

yoboijarvis:

0.9??

AZ:

Exactly!! Now remember, how we said that we're increasing her salary by 10%? that means you just add that 0.9 to the original 9 so her new salary is basically 9 + 0.9 = ??

yoboijarvis:

9 . 9

AZ:

You got it! Another way to do it since we're INCREASING by 10% we know that 9 is the whole 100% and we're increasing by 10% we can just multiply 9 by (100% + 10%) or by 110% so 110% = 110/100 and multiplying it with 9 would get you the same answer

yoboijarvis:

Okay, thank you very much AZ.

AZ:

You're welcome!